A parabola given by x = ay2 + by + c passes through points (1,2) , (-3,2) and (-2,4) .



1. Find the equation ?



2. Determine whether the line x + 2y + 3 = 0 is tangent to the parabola in (1)



3. Show that each point P(x,y) on the parabola is equidistant from (-2,2) and the line x= -4

1
Expert's answer
2022-04-07T09:28:05-0400

The parabola has the form: "x=ay^2+by+c". Substitute points "(1,2), (-3,2)" and "(-2,4)" into equation of parabola: "1=4a+2b+c," "-3=4a+2b+c" and "-2=16a+4b+c". We receive: "-3=1" from the first and second equations. Thus, there is a mistake in the formulation of the problem. The points are not chosen correctly.